# Real-Time Sampling-Based Safe Motion Planning for Robotic Manipulators in Dynamic Environments

### v1

@@ IV-A Dynamic Expanded Bubbles and Burs
First, in order to use (4), a local path $\overline{\boldsymbol{q}\boldsymbol{y}}$, which is a straight-line segment (i.e., spine [54]), must be time-parameterized since $\boldsymbol{y}=\boldsymbol{y}(\widetilde{\boldsymbol{q}},t)$, obtained for the specific value of $t$, is a time-dependent configuration, where $\boldsymbol{y}(\widetilde{\boldsymbol{q}},t)$ represents a family of all possible local trajectories $\boldsymbol{\pi}(t)=\boldsymbol{\pi}[\boldsymbol{q},\boldsymbol{y}(\widetilde{\boldsymbol{q}},t^{*})]$. Computing the intersection of the specific trajectory $\boldsymbol{\pi}(t)$ with the border of $\mathcal{DEB}$ is obtained by solving the system of equations $\sum_{j=1}^{i}r_{i,j}\left|\pi_{i}(t)-q_{i}\right|+v_{obs}t=d_{i}$, $\forall i\in\{1,2,\dots,n\}$. Multiple solutions may exist due to the nonlinear nature of $\boldsymbol{\pi}(t)$ (e.g., polynomial). Clearly, the smallest value of $t=t^{*}$, where $t^{*}\in(0,\frac{d_{c}}{v_{obs}})$, should be computed. Since it could be relatively computationally expensive, we do not seek the exact value of $t^{*}$. Instead, we resort to a convenient workaround described in the sequel within Alg. 2. The procedure generates so-called dynamic bur $DBur$ using the following steps:
Definition IV.4.
Figure 6: Process of computing $DBur_{0}$ (solid black lines)
@@ 23             $DBur\leftarrow[DBur,\,\overline{\boldsymbol{q}_{0}\boldsymbol{q}_{k}}]$
Figure 7: (a) Process of computing dynamic generalized bur (solid yellow lines); (b)/(c)/(d) The nearest points, $\color[rgb]{1,0,0}\boldsymbol{R}(t)\equiv\boldsymbol{R}_{i,j}(t)$ and $\color[rgb]{1,0,0}\boldsymbol{O}(t)\equiv\boldsymbol{O}_{i,j}(t)$, from $i$-th robot’s link and $\mathcal{WO}_{j}$, respectively, with the corresponding plane $\color[rgb]{0,0,1}\mathcal{P}(t)\equiv\mathcal{P}_{i,j}(t)$ and the distance $d(t)=d_{i,j}(t)$ after exceeding the time $t\in\{t_{0},t_{3},t_{7}\}$.
IV-B Chaining Dynamic Expanded Bubbles and Burs
We are particularly interested in further exploration of $\mathcal{C}_{free}$ after reaching the border of $\mathcal{DEB}$ (e.g., $\boldsymbol{q}_{6}$ in Fig. 6) since spines $\overline{\boldsymbol{q}_{0}\boldsymbol{q}_{k}}\notin DBur_{0}$ do not have to be in a collision (e.g., spines $\overline{\boldsymbol{q}_{0}\boldsymbol{q}_{7}}$, $\overline{\boldsymbol{q}_{0}\boldsymbol{q}_{8}}$ and $\overline{\boldsymbol{q}_{0}\boldsymbol{q}_{9}}$ from Fig. 6). To this end, we propose a simple and fast approach within Alg. 3 for computing so-called dynamic generalized bur $DGBur$. The proposed procedure involves the following steps:
Step 1 (lines 33): First, $DBur_{k}$ is generated by the $\mathtt{computeDBur}$ function (provided by Alg. 2) using intermediate nodes from the trajectory $\boldsymbol{\pi}[\boldsymbol{q}_{k},\boldsymbol{q}_{f}]$, $\forall k\in\{m,\dots,N-1\}$, for some $m\in\{0,1,\dots,N-1\}$, (e.g., yellow spines from $\boldsymbol{q}_{0}$ towards $\boldsymbol{q}_{1},\boldsymbol{q}_{2},\boldsymbol{q}_{3}\in\mathcal{DEB}_{0}$, from $\boldsymbol{q}_{3}$ towards $\boldsymbol{q}_{4},\dots,\boldsymbol{q}_{7}\in\mathcal{DEB}_{3}$, and from $\boldsymbol{q}_{7}$ towards $\boldsymbol{q}_{8},\dots,\boldsymbol{q}_{11}\in\mathcal{DEB}_{7}$ in Fig. 7 (a)). Hence, if $DBur_{k}$ exists, the subsequent steps can be executed. Otherwise, the procedure returns $DGBur$ computed so far.
Step 4 (line 3): Since the robot has moved from $\boldsymbol{q}_{k}$ to $\boldsymbol{q}_{m}$, new underestimates of $\boldsymbol{d}$ can be easily acquired (see [51]) by the $\mathtt{getDistancesToPlanes}$ function as distances to the new planes $\boldsymbol{\mathcal{P}}(t_{m})$, when the robot assumes $\boldsymbol{q}_{m}$ (e.g., $\boldsymbol{q}_{3}$ and $\boldsymbol{q}_{7}$ in Fig. 7 (c)–(d)). Interestingly, the proposed way of computing distances may even increase the size of $\mathcal{DEB}_{m}$ in case the $i$-th link moves away from $\mathcal{WO}_{j}$ providing $d_{i}(t_{m})>d_{i}(t_{k})$.
Step 5 (line 33): All spines from the newly computed bur $DBur_{k}$ are concatenated within $DGBur$. The size of a chain of connected $\mathcal{DEB}$s (it is the same as the number of generated corresponding dynamic burs) is tracked via the local variable $chain\_size$. It also denotes the number of extensions/layers of the dynamic generalized bur (e.g., Fig. 7 (a) depicts 3 layers in total).
The logic from all steps 1–5 can be repeated for the new $k\leftarrow m$ until all intermediate nodes from $\boldsymbol{\pi}[\boldsymbol{q}_{0},\boldsymbol{q}_{f}]$ turn out to be collision-free (i.e., while $\boldsymbol{q}_{k}\neq\boldsymbol{q}_{f}$), or until the maximal number of layers $K$ is reached, or until $DBur_{k}=\emptyset$.
@@ IV-B Chaining Dynamic Expanded Bubbles and Burs
11      if $DBur_{k}\neq\emptyset$ then
@@ 25            $k\leftarrow m$
The set of collision-free spines $\overline{\boldsymbol{q}_{k}\boldsymbol{q}_{i}}\in\mathcal{DEB}(\widetilde{\boldsymbol{q}}_{k},\boldsymbol{d}(t_{k}),v_{obs})$, $\forall i\in\{k+1,\dots,m\}$, for some $k\in\{0,1,\dots,$ $N-1\}$ and $m\in\{k+1,\dots,$ $N\}$, represents the structure we refer to as dynamic generalized bur.
@@
Figure 8: The set of collision-free trajectories in $\{q_{1},q_{2}\}$-plane generated, using the notion of $DGBur$, from the configurations depicted in Fig. 5 (left)
@@ IV-C Generating trajectories
The process of generating splines is a part of the $\mathtt{updateCurrState}$ routine from Alg. 1 depending on the parameter $safe\_on$ indicating whether the spline is guaranteedly safe for the environment. Our approach focuses on computing the spline with its shortest possible duration time $t_{f}$ while satisfying the set of constraints $\mathcal{K}$ (i.e., maximal joint velocity, acceleration, and jerk). The initial position $\boldsymbol{\pi}_{curr}(t_{0})=\boldsymbol{q}_{curr}$, velocity $\dot{\boldsymbol{\pi}}_{curr}(t_{0})$, and acceleration $\ddot{\boldsymbol{\pi}}_{curr}(t_{0})$ are inherited from the spline computed in the previous iteration. The final position $\boldsymbol{\pi}_{curr}(t_{f})=\boldsymbol{q}_{next}$ remains fixed. In case $safe\_on$ is false, the final velocity $\dot{\boldsymbol{\pi}}_{curr}(t_{f})$ is estimated using two configurations, $\boldsymbol{q}_{curr}$ and $\boldsymbol{q}_{next}$, in order to achieve a minimal value for $t_{f}$ to account for $\mathcal{K}$. Otherwise, when $safe\_on$ is true, $\dot{\boldsymbol{\pi}}_{curr}(t_{f})=\boldsymbol{0}$. For simplicity, the final acceleration is always set to zero, i.e., $\ddot{\boldsymbol{\pi}}_{curr}(t_{f})=\boldsymbol{0}$. Using these six equations suffices to compute the quintic spline $\boldsymbol{\pi}_{curr}(t)$.
@@ IV-D Application of Dynamic Expanded Bubbles and Burs
In case $\boldsymbol{\pi}_{comp}(t)$ is not a safe spline, then $\boldsymbol{q}_{next}$ should be changed and a new spline $\boldsymbol{\pi}_{comp}(t)$ acquired. Finally, the safety of that spline can be checked. We propose using bisection method as shown in Fig. 9 over the horizon spine $\overline{\boldsymbol{q}_{curr}\boldsymbol{q}_{next}^{{}_{(0)}}}$ (dotted black line) in order to find a new convenient node $\boldsymbol{q}_{next}$ that will yield a safe spline. For instance, assume $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{stop}^{(0)}]$ is unsafe. Then, a new next configuration is computed as $\boldsymbol{q}_{next}^{(1)}=(\boldsymbol{q}_{curr}+\boldsymbol{q}_{next}^{(0)})/2$, and $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{stop}^{(1)}]$ is checked for safety. If it is safe, the process might terminate. However, in search for a longer spline, a more distant node is computed as $\boldsymbol{q}_{next}^{(2)}=(\boldsymbol{q}_{next}^{(1)}+\boldsymbol{q}_{next}^{(0)})/2$, and $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{stop}^{(2)}]$ is checked for safety as illustrated by Fig. 7 (a) in detail. Such a bisection method can run until the maximal number of iterations exceeds. If this approach never finds a safe spline, there is always an ultimate solution – immediately execute an emergency stopping from $\boldsymbol{q}_{curr}$, thus safely stopping the robot at the configuration $\boldsymbol{q}_{stop}^{\prime}$ (e.g., dotted red line in Fig. 7 (a)). The stopping trajectory $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{stop}^{\prime}]$ will also be safe since it is a part of the safe spline that was computed in the previous algorithm iteration.
Figure 9: Using bisection method to find a safe spline
@@
Theorem 1 (Guaranteed safe motion of the robot in dynamic environments).
If the configuration $\boldsymbol{q}_{new}$ is safe, the robot will follow the safe spline $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{new}]$ during the time $t\in[t_{0},t_{new}]$, thus moving from $\boldsymbol{q}_{curr}$ to $\boldsymbol{q}_{new}$ satisfying the set of constraints $\mathcal{K}$. Otherwise, the robot will immediately execute stopping from $\boldsymbol{q}_{curr}$ to $\boldsymbol{q}_{stop}^{\prime}$ by following the safe spline $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{stop}^{\prime}]$. Therefore, a solution satisfying $\mathcal{K}$ always exists leading the robot safely within the boundaries of the chain of connected dynamic expanded bubbles.
@@
The robot will automatically scale its velocity depending on the size of obtained $\mathcal{DEB}$s that further depends on $d_{i}(t_{k})$, as well as on the direction of the robot’s links motion. For smaller values of $d_{i}(t_{k})$, and particularly when the robot’s $i$-th link approaches $\mathcal{WO}_{j}$, $\mathcal{DEB}$s become smaller. Consequently, corresponding safe splines get shorter. Since the splines are computed with the shortest possible $t_{f}$, and regarding $\dot{\boldsymbol{q}}_{next}=\boldsymbol{0}$ and $\ddot{\boldsymbol{q}}_{next}=\boldsymbol{0}$, the maximal velocity on $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{next}]$ reduces if $\boldsymbol{q}_{next}$ is closer to $\boldsymbol{q}_{curr}$ (e.g., the maximal velocity on $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{{}_{next}}^{{}_{(2)}}]$ is smaller than the one on $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{{}_{next}}^{{}_{(0)}}]$ when considering Fig. 9).

### v2

@@ IV-A Dynamic Expanded Bubbles and Burs
Figure 6: Process of computing dynamic bur (solid black lines).
Generally, any motion within $\mathcal{DEB}$ means that no point on the robot $\mathcal{R}_{i}$ will move more than $d_{i}-v_{obs}t$ for a given time $t$, thus no collision can occur. Such a defined bubble differs from the one from [57] in terms of reducing the initial value of each $d_{i}$ for $v_{obs}t$. The rationale is that each $\mathcal{WO}_{j}$, may cover the distance of $v_{obs}t$ in the worst-case scenario, i.e., it may approach the robot’s $i$-th link at its highest speed $v_{obs}$. This observation is equivalent to considering the plane $\mathcal{P}_{i,j}$ which artificially enlarges $\mathcal{WO}_{j}$ by following a velocity vector $\boldsymbol{v}_{\mathcal{P}_{i,j}}=\frac{\boldsymbol{R}_{i,j}-\boldsymbol{O}_{i,j}}{\|\boldsymbol{R}_{i,j}-\boldsymbol{O}_{i,j}\|}v_{obs}$ (as depicted in Fig. 7 (b, c, d)). It is worth mentioning that we assume unpredictable motion of obstacles without monitoring or predicting their directions of motion. Otherwise, it would reduce the magnitude of the approaching vector $\boldsymbol{v}_{\mathcal{P}_{i,j}}$ from $v_{obs}$ to $v_{obs}\cdot\cos\varangle(\boldsymbol{v}_{\mathcal{P}_{i,j}},\boldsymbol{v}_{j})$, where $\boldsymbol{v}_{j}$ represents a monitored $j$-th obstacle velocity vector.
Fig. 5 illustrates the concept of dynamic bubbles. The workspace of a planar 2-DoF robot in four different configurations with two box-shaped obstacles is shown in Fig. 5 (left). Changing $v_{obs}$ from 0 to $6\,\mathrm{[\frac{m}{s}]}$ affects the $\mathcal{DEB}$ with a root in $\boldsymbol{q}_{0}=\left[-\frac{\pi}{4},0\right]^{T}$ (blue configuration) is revealed by Fig. 5 (right). Note that the case for $v_{obs}=0$ corresponds with the expanded bubble from [57]. For higher velocities, $\mathcal{DEB}$s get smaller, and their shape deviates from the diamond-like one. The initial velocity and acceleration used for time parameterization of trajectories are set to zero in all cases.
First, in order to use (5), a local path $\overline{\boldsymbol{q}\boldsymbol{y}}$, which is a straight-line segment (i.e., spine [55]), must be time-parameterized since $\boldsymbol{y}=\boldsymbol{y}(\widetilde{\boldsymbol{q}},t)$, obtained for the specific value of $t$, is a time-dependent configuration, where $\boldsymbol{y}(\widetilde{\boldsymbol{q}},t)$ represents a family of all possible local trajectories $\boldsymbol{\pi}(t)=\boldsymbol{\pi}[\boldsymbol{q},\boldsymbol{y}(\widetilde{\boldsymbol{q}},t^{*})]$. Computing the intersection of the specific trajectory $\boldsymbol{\pi}(t)$ with the border of $\mathcal{DEB}$ is obtained by solving the system of equations $\sum_{j=1}^{i}r_{i,j}\left|\pi_{j}(t)-q_{j}\right|+v_{obs}t=d_{i}$, $\forall i\in\{1,2,\dots,n\}$. Multiple solutions may exist due to the nonlinear nature of $\boldsymbol{\pi}(t)$ (e.g., polynomial). Clearly, the smallest value of $t=t^{*}$, where $t^{*}\in(0,\frac{d_{c}}{v_{obs}})$, should be computed. Since it could be relatively computationally expensive, we do not seek the exact value of $t^{*}$. Instead, we resort to a convenient workaround described in the sequel within Alg. 2. The procedure generates a so-called dynamic bur $DBur$ using the following steps:
Figure 7: (a) Process of computing dynamic generalized bur (solid yellow lines); (b,c,d) The nearest points, $\color[rgb]{1,0,0}\boldsymbol{R}(t)\equiv\boldsymbol{R}_{i,j}(t)$ and $\color[rgb]{1,0,0}\boldsymbol{O}(t)\equiv\boldsymbol{O}_{i,j}(t)$, from $i$-th robot’s link and $\mathcal{WO}_{j}$, with the corresponding plane $\color[rgb]{0,0,1}\mathcal{P}(t)\equiv\mathcal{P}_{i,j}(t)$ and the distance $d(t)\equiv d_{i,j}(t)$ after $t\in\{t_{0},t_{3},t_{7}\}$.
@@ 23     $DBur\leftarrow[DBur,\,\overline{\boldsymbol{q}_{0}\boldsymbol{q}_{k}}]$
Definition IV.4.
@@
11 if $DBur_{k}\neq\varnothing$ then
@@ 25    $k\leftarrow m$
IV-B Chaining Dynamic Expanded Bubbles and Burs
We are particularly interested in further exploration of $\mathcal{C}_{free}$ after reaching the border of $\mathcal{DEB}$ (e.g., $\boldsymbol{q}_{6}$ in Fig. 6) since spines $\overline{\boldsymbol{q}_{0}\boldsymbol{q}_{k}}\notin DBur_{0}$ do not have to be in a collision (e.g., spines $\overline{\boldsymbol{q}_{0}\boldsymbol{q}_{7}}$, $\overline{\boldsymbol{q}_{0}\boldsymbol{q}_{8}}$ and $\overline{\boldsymbol{q}_{0}\boldsymbol{q}_{9}}$ from Fig. 6). To this end, we propose a simple and fast approach within Alg. 3 for computing a so-called dynamic generalized bur $DGBur$. The proposed procedure involves the following steps:
Step 1 (lines 33): First, $DBur_{k}$ is generated by the $\mathtt{computeDBur}$ function (provided by Alg. 2) using intermediate nodes from the trajectory $\boldsymbol{\pi}[\boldsymbol{q}_{k},\boldsymbol{q}_{f}]$, $\forall k\in\{m,\dots,N-1\}$, for some $m\in\{0,1,\dots,N-1\}$, (e.g., yellow spines from $\boldsymbol{q}_{0}$ towards $\boldsymbol{q}_{1},\boldsymbol{q}_{2},\boldsymbol{q}_{3}\in\mathcal{DEB}_{0}$, from $\boldsymbol{q}_{3}$ towards $\boldsymbol{q}_{4},\dots,\boldsymbol{q}_{7}\in\mathcal{DEB}_{3}$, and from $\boldsymbol{q}_{7}$ towards $\boldsymbol{q}_{8},\dots,\boldsymbol{q}_{11}\in\mathcal{DEB}_{7}$ in Fig. 7 (a)). If $DBur_{k}$ exists, the subsequent steps can be executed. Otherwise, the procedure returns $DGBur$ computed so far.
Step 4 (line 3): Since the robot has moved from $\boldsymbol{q}_{k}$ to $\boldsymbol{q}_{m}$, new underestimates of $\boldsymbol{d}$ can be easily acquired (see [51]) by the $\mathtt{getDistancesToPlanes}$ function as distances to the new planes $\boldsymbol{\mathcal{P}}(t_{m})$, when the robot assumes $\boldsymbol{q}_{m}$ (e.g., $\boldsymbol{q}_{3}$ and $\boldsymbol{q}_{7}$ in Fig. 7 (c, d)). Interestingly, the proposed way of computing distances may even increase the size of $\mathcal{DEB}_{m}$ in case the $i$-th link moves away from $\mathcal{WO}_{j}$ providing $d_{i}(t_{m})>d_{i}(t_{k})$.
Step 5 (line 33): All spines from the newly computed bur $DBur_{k}$ are concatenated within $DGBur$. The size of a chain of connected $\mathcal{DEB}$s (it is the same as the number of generated corresponding dynamic burs) is tracked via the local variable chain_sizechain\_size. It also denotes the number of extensions/layers of the dynamic generalized bur (e.g., Fig. 7 (a) depicts 3 layers in total).
The logic from all steps 1–5 can be repeated for the new $k\leftarrow m$ until all intermediate nodes from $\boldsymbol{\pi}[\boldsymbol{q}_{0},\boldsymbol{q}_{f}]$ turn out to be collision-free (i.e., while $\boldsymbol{q}_{k}\neq\boldsymbol{q}_{f}$), or until the maximal number of layers $K$ is reached, or until $DBur_{k}=\varnothing$.
@@ IV-B Chaining Dynamic Expanded Bubbles and Burs
The set of collision-free spines $\overline{\boldsymbol{q}_{k}\boldsymbol{q}_{i}}\in\mathcal{DEB}(\widetilde{\boldsymbol{q}}_{k},\boldsymbol{d}(t_{k}),v_{obs})$, $\forall i\in\{k+1,\dots,m\}$, for some $k\in\{0,1,\dots,$ N1}N-1\} and $m\in\{k+1,\dots,$ N}N\}, represents the structure we refer to as dynamic generalized bur.
@@
Figure 8: The set of collision-free trajectories in $\{q_{1},q_{2}\}$-plane, generated using $DGBur$, from the configurations depicted in Fig. 5 (left).
@@ IV-C Generating Trajectories
The process of generating splines is a part of the $\mathtt{updateCurrState}$ routine from Alg. 1 depending on the parameter safe_onsafe\_on indicating whether the spline is guaranteedly safe for the environment. Our approach focuses on computing the spline with its shortest possible duration time $t_{f}$ while satisfying the set of constraints $\mathcal{K}$ (i.e., maximal joint velocity, acceleration, and jerk). The initial position $\boldsymbol{\pi}_{curr}(t_{0})=\boldsymbol{q}_{curr}$, velocity $\dot{\boldsymbol{\pi}}_{curr}(t_{0})$, and acceleration $\ddot{\boldsymbol{\pi}}_{curr}(t_{0})$ are inherited from the spline computed in the previous iteration. The final position $\boldsymbol{\pi}_{curr}(t_{f})=\boldsymbol{q}_{next}$ remains fixed. In case safe_onsafe\_on is false, the final velocity $\dot{\boldsymbol{\pi}}_{curr}(t_{f})$ is estimated using two configurations, $\boldsymbol{q}_{curr}$ and $\boldsymbol{q}_{next}$, in order to achieve a minimal value for $t_{f}$ to account for $\mathcal{K}$. Otherwise, when safe_onsafe\_on is true, $\dot{\boldsymbol{\pi}}_{curr}(t_{f})=\boldsymbol{0}$. For simplicity, the final acceleration is always set to zero, i.e., $\ddot{\boldsymbol{\pi}}_{curr}(t_{f})=\boldsymbol{0}$. Using these six equations suffices to compute the quintic spline $\boldsymbol{\pi}_{curr}(t)$.
@@ IV-D Application of Dynamic Expanded Bubbles and Burs
In case $\boldsymbol{\pi}_{comp}(t)$ is possibly unsafe, then $\boldsymbol{q}_{next}$ should be changed and a new spline $\boldsymbol{\pi}_{comp}(t)$ acquired. Finally, the safety of that spline can be checked. We propose using bisection method as shown in Fig. 9 over the horizon spine $\overline{\boldsymbol{q}_{curr}\boldsymbol{q}_{next}^{{}_{(0)}}}$ (dotted black line) in order to find a new convenient node $\boldsymbol{q}_{next}$ that will yield a safe spline. For instance, assume $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{stop}^{(0)}]$ is unsafe. Then, a new next configuration is computed as $\boldsymbol{q}_{next}^{(1)}=(\boldsymbol{q}_{curr}+\boldsymbol{q}_{next}^{(0)})/2$, and $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{stop}^{(1)}]$ is checked for safety. If it is safe, the process might terminate. However, in search for a longer spline, a more distant node is computed as $\boldsymbol{q}_{next}^{(2)}=(\boldsymbol{q}_{next}^{(1)}+\boldsymbol{q}_{next}^{(0)})/2$, and $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{stop}^{(2)}]$ is checked for safety as illustrated by Fig. 7 (a) in detail. The described problem can be formulated as:
Problem IV.1 (Computing a safe spline).
For a given desired next configuration $\boldsymbol{q}_{next}$, we aim at computing the longest possible spline from the current configuration $\boldsymbol{q}_{curr}$ towards $\boldsymbol{q}_{next}$ which is guaranteedly safe, and which has minimal final time $t_{f}$ while respecting the set of constraints $\mathcal{K}$.
Such a bisection method can run until the maximal number of iterations exceeds. If this approach never finds a safe spline, there is always an ultimate solution – immediately execute an emergency stopping from $\boldsymbol{q}_{curr}$, thus safely stopping the robot at the configuration $\boldsymbol{q}_{stop}^{\prime}$ (e.g., dotted red line in Fig. 7 (a)). The stopping trajectory $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{stop}^{\prime}]$ will also be safe since it is a part of the safe spline that was computed in the previous algorithm iteration.
Figure 9: Using bisection method to find a safe spline.
@@
Theorem 1 (Guaranteed safe motion of the robot in dynamic environments under bounded obstacle velocity).
If a configuration $\boldsymbol{q}_{new}$ is safe and the maximal obstacle velocity is limited to $v_{obs}$, the robot will follow the safe spline $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{new}]$ during the time $t\in[t_{0},t_{new}]$, thus moving from the current configuration $\boldsymbol{q}_{curr}$ to a new configuration $\boldsymbol{q}_{new}$ satisfying the set of kinematic constraints $\mathcal{K}$. Otherwise, the robot will immediately execute stopping from $\boldsymbol{q}_{curr}$ to another safe configuration $\boldsymbol{q}_{stop}^{\prime}$ by following the safe spline $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{stop}^{\prime}]$. Therefore, a solution satisfying $\mathcal{K}$ always exists leading the robot safely within the boundaries of the chain of connected dynamic expanded bubbles.
@@
The robot will automatically scale its velocity depending on the size of obtained $\mathcal{DEB}$s that further depends on $d_{i}(t_{k})$, as well as on the direction of the robot’s links motion. For smaller values of $d_{i}(t_{k})$, and particularly when the robot’s $i$-th link approaches $\mathcal{WO}_{j}$, $\mathcal{DEB}$s become smaller. Consequently, corresponding safe splines get shorter. Regarding Problem IV.1 as well as $\dot{\boldsymbol{q}}_{next}=\boldsymbol{0}$ and $\ddot{\boldsymbol{q}}_{next}=\boldsymbol{0}$, the obtained maximal velocity on $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{next}]$ may decrease if $\boldsymbol{q}_{next}$ is closer to $\boldsymbol{q}_{curr}$ (e.g., the obtained maximal velocity on $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{next}^{{}_{(2)}}]$ is smaller or equal to the one on $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{next}^{{}_{(0)}}]$ when considering Fig. 9).